OpenStudy (bbb911):

Multiply to the lowest term 5/8 X 18/35 X 4/9

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

\[\frac{ 360 }{ 2520 } = \frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

its 1/5

OpenStudy (bbb911):

wait i am confused

OpenStudy (bbb911):

how did you get 1/5

OpenStudy (anonymous):

multiply them all you would get 0.2 in fraction mode its 1/5

OpenStudy (bbb911):

i am just trying to stay away from decimals and just simplify them till it gets to a lowest term

OpenStudy (anonymous):

5 x 18 x 4 = 360 8 x 35 x 9 = 2520 So... \[\frac{ 360 }{ 2520 } = \frac{ 1 }{ 7 }\]

OpenStudy (bbb911):

thank you so mcuh:)

OpenStudy (anonymous):

your right my bad